OpenStudy (anonymous):

What are the solutions of 3x2 = 18x? x = 0, x = 6 x = 0, x = −6 x = 3, x = 6 x = 3, x = −6

OpenStudy (lgbasallote):

subtract 18x from both sides first..what do you get?

OpenStudy (anonymous):

okay one sec

OpenStudy (anonymous):

okay wait im confused soo do the exponents first right..?

Directrix (directrix):

3x^2 = 18x 3x^2 - 18x = 0 Next, factor out 3x from both expressions on the left.

OpenStudy (anonymous):

soo is it x=3 and x=-6..?

Directrix (directrix):

Factor first. 3x^2 - 18x = 0 3x(x - 6) = 0 Then, use the Zero Product Property which states that if two factors multiply to zero, then one or both factors are equal to zero. So, 3x = 0 or x - 6 = 0 x = 0 or x = 6 --> Answer

OpenStudy (anonymous):

ohh dang i was wrong..lol i suck and thank you soo much !

Directrix (directrix):

Glad to help. You just skipped a step.